# WaveGuard: Robust Deepfake Detection and Source Tracing via Dual-Tree Complex Wavelet and Graph Neural Networks

### v3

@@ IV-C Robustness Test
SaltPepper 0.0078% 0.0178% 0.0123% 0.0083% 0.0156%
GaussianNoise 0.0149% 0.0128% 0.7410% 0.0132% 0.0327%
Average 0.0032% 0.0045% 0.0999% 0.0030% 0.0065%
SimSwap 45.4832% 46.3146% 46.7812% 46.1530% 46.7845%
GANimation 45.4851% 46.3144% 46.7856% 46.1534% 46.7849%
StarGAN (Male) 45.4835% 46.3132% 46.7837% 46.1227% 46.7828%
StarGAN (Young) 45.4842% 46.3187% 46.7836% 46.1510% 46.7828%
StarGAN (BlackHair) 45.4811% 46.3131% 46.7842% 46.1574% 46.7845%
StarGAN (BlondHair) 45.4896% 46.3145% 46.7813% 46.1533% 46.7811%
StarGAN (BrownHair) 45.4842% 46.3149% 46.7811% 46.2061% 46.7828%
Average 45.4844% 46.3148% 46.7830% 46.1567% 46.7833%
TABLE II quantitatively compares the BER performance of different methods under common distortion and deepfake forgery, covering 128×128 and 256×256 resolutions. Experimental results show that our method achieves the lowest BER in all test scenarios, showing excellent robustness. Under ordinary distortion, the BER of Tracer and Detector are both close to 0%, which ensures that the watermark can be recovered with high precision under all kinds of interference. However, the BER of the contrast method is obviously higher under various common distortion conditions, which indicates that its robustness is weak. Under the deepfake forgery, the BER of Tracer is still close to 0%, while the BER of semi-robust Detector is maintained close to 50%, which effectively distinguishes the true and forged contents. Although our proposed semi-robust Detector exhibits a slightly lower BER than SepMark under deepfake attacks, the primary objective of the Detector is to distinguish between authentic and forged content. As long as a significant BER gap exists between common distortions and deepfake forgeries, effective detection can be achieved. Our method yields a substantially lower BER under common distortions, thereby enlarging the separation margin between real and fake samples and significantly enhancing discriminability. However, the existing methods have large fluctuations in BER and lack of detection reliability.
To evaluate the cross-dataset generalization ability, TABLE III reports the BER performance of our method on the LFW dataset under both common and malicious distortions at two resolutions. The values in bold represent the absolute difference in BER between LFW and the original dataset (CelebA-HQ), with arrows indicating the direction of change—$\uparrow$ for increased BER and $\downarrow$ for decreased BER on LFW. Despite the domain shift, our method maintains low BER in tracing and preserves the 50% boundary in detection across most distortion types. This suggests that our framework is not only robust under in-domain perturbations but also generalizes well to unseen data from different distributions, demonstrating strong transferability and practical applicability.
This robustness is mainly due to the frequency domain embedding strategy. Compared with the spatial domain method, the frequency domain watermark is more stable under common distortion, ensuring the integrity of the watermark information. At the same time, the design of joint sub-band enhances the discrimination ability of semi-robust Detector under deepfake forgery, so that it can stably distinguish real and forged content.
IV-D Ablation Study
@@
In order to evaluate the contribution of Joint Sub-bands, Attention Module, Frequency Domain Embedding, and GNN in this method, we conducted ablation experiments on both Tracer and Detector. The experiments were conducted on images with a resolution of 128×128. (a) Adjust the sub-bands usage of Tracer and Detector: In Tracer, watermark embedding was reduced from 1,3,4,6 sub-bands to 1,3, while in Detector, the detection was expanded from 1,3 sub-bands to 1,3,4,6. This allows us to evaluate the influence of joint sub-band information redundancy on watermark recovery and deepfake detection. (b) Remove the attention mechanism in the process of watermark embedding to analyze its influence on visual quality. (c) When the DT-CWT frequency domain embedding is removed, the watermark is directly embedded in the U channel. (d) Removing SC-GNN to evaluate its impact on the visual quality of the watermarked image. TABLE V and V present the results of the ablation study.
@@
It can be found that: (a) From Case 1, the joint sub-band significantly affects the robustness and detection ability of the watermark. When Tracer only extracts the watermark in 1 and 3 sub-bands, the BER increases obviously, indicating that reducing sub-bands weakens the stability of the watermark. However, when the Detector is extended to 1, 3, 4 and 6 sub-bands, the BER decreases, and the extra sub-bands enhance the redundancy of the watermark, making its detection more stable under common distortion and deepfake forgery. (b) From Case 2, it can be seen that the PSNR and SSIM of the watermark image decrease after the attention mechanism is removed, which indicates that the attention mechanism has optimized the watermark embedding. Although BER has not changed much, the invisibility and visual quality of the watermark have obviously decreased, which verifies the importance of attention mechanism in optimizing visual quality. (c) From Case 3, it can be seen that the robustness of the watermark is significantly reduced after the frequency domain embedding is removed. Although PSNR and SSIM remain at a high level, BER increases significantly under common distortion and deepfake forgery. In addition,the perfect reconstruction, directional selectivity and approximate translation invariance of DT-CWT make frequency domain embedding more resistant to distortion and forgery, which verifies its key role in improving watermark robustness. (d) From Case 4, it can be seen that after removing SC-GNN, the PSNR and SSIM of the watermark image decrease, although the BER changes little, indicating that SC-GNN plays an important role in optimizing the visual quality of the watermark image. SC-GNN improves the invisibility of the watermark and reduces perceptual artifacts by extracting the structural features of the image, which verifies the key role of SC-GNN in optimizing the visual quality of the watermark.
@@ V CONCLUSION
In this paper, a watermarking framework based on frequency domain embedding and Structural Consistency Graph Neural Network (SC-GNN) is proposed. The robustness is improved by Dual-Tree Complex Wavelet Transform (DT-CWT) transform, and the detection ability of deepfake is enhanced by combining sub-band information. In addition, the framework leverages dense connectivity to enhance feature reuse and robustness. SC-GNN further optimizes the visual quality of watermark embedding, reduces perceptual artifacts, and improves the invisibility of watermark by attention mechanism. Experimental results show that this method can stably extract watermark and accurately detect tampering in various distorted and forged scenes. Future work will further improve its generalization ability and optimize its adaptability to complex distortion and unknown deepfake technology.

### v4

@@ IV-C Robustness Test
SaltPepper 0.0277% 0.0298% 0.0308% 0.0081% 0.0211%
GaussianNoise 0.0149% 0.0121% 0.7391% 0.0209% 0.0361%
Average 0.0072% 0.0071% 0.1189% 0.0056% 0.0089%
FSRT 45.4874% 46.3172% 46.7846% 46.1524% 46.7852%
SimSwap 45.4821% 46.3141% 46.7871% 46.1574% 46.7815%
CSCS 45.4832% 46.3164% 46.7873% 46.1524% 46.7879%
HifiFace 45.4852% 46.3119% 46.7877% 46.1499% 46.7872%
Uniface 45.4840% 46.3177% 46.7857% 46.1577% 46.7871%
GANimation 45.4851% 46.3134% 46.7856% 46.1525% 46.7873%
StarGAN (Male) 45.4834% 46.3122% 46.7836% 46.1531% 46.7869%
StarGAN (Young) 45.4822% 46.3188% 46.7856% 46.1510% 46.7875%
StarGAN (BlackHair) 45.4831% 46.3139% 46.7862% 46.1584% 46.7876%
StarGAN (BlondHair) 45.4816% 46.3147% 46.7843% 46.1545% 46.7876%
StarGAN (BrownHair) 45.4843% 46.3147% 46.7851% 46.1547% 46.7872%
Average 45.4738% 46.3162% 46.7857% 46.1586% 46.7848%
[…]
This robustness is mainly due to the frequency-domain embedding strategy, which plays a central role in ensuring generalization across datasets. By distributing watermark information across multiple directional sub-bands, the DT-CWT enhances redundancy and stability under various distortions and data distributions. In addition, the SC-GNN module provides structural modeling that helps the system capture cross-domain spatial consistencies, further improving robustness against semantic and appearance variations. Finally, the feature reuse mechanism is introduced to improve the flexibility and robustness of watermark embedding by optimizing the feature transfer process.
Fig. 8: Sensitivity analysis of watermark robustness under diverse deepfake distortions at varying resolutions.
TABLE VII: Computational complexity of WaveGuard: parameters, FLOPs, and inference time on an NVIDIA RTX 4090 GPU.
Input Resolution Params (M) FLOPs (G) Time (ms/image)
128 $\times$ 128 15.5 3.38 74.0
256 $\times$ 256 15.5 13.5 70.9
IV-D Complexity Analysis
To comprehensively evaluate the computational feasibility of WaveGuard, we report its parameter size, FLOPs, and inference speed on an NVIDIA RTX 4090 GPU, as shown in TABLE VII. The model contains approximately 15.5M trainable parameters, reflecting moderate structural complexity. In terms of computational cost, WaveGuard requires about 13.5 GFLOPs to process a $256\times 256$ image, with an average inference time of 70.9 ms per image. When the input resolution is reduced to $128\times 128$, the computational cost drops significantly to about 3.38 GFLOPs (roughly one quarter of the $256\times 256$ case due to the quadratic reduction in pixel count), while the inference time remains similar at around 74.0 ms per image owing to GPU parallelization. These results indicate that WaveGuard achieves a well-balanced trade-off between robustness, traceability, and computational efficiency, making its complexity acceptable for practical deepfake detection and source tracing scenarios.
IV-E Robustness Sensitivity Analysis
To analyze the robustness and sensitivity of watermark embedding under different deepfake generation types and image resolutions (128×128 and 256×256), we present the evaluation results in Fig. 8. Among them, the BER of SimSwap and FSRT increases significantly at low resolution, which indicates that their operations based on semantic reconstruction and whole face replacement have the most serious damage to the watermark. However, forgery methods such as CSCS, HifiFace and GANimation have limited changes to the image structure, so the BER is always kept at a low level. The average BER of StarGAN series is also low, indicating that the interference ability of attribute-level forgery to watermark is relatively weak. On the whole, the robustness of 256×256 resolution is significantly better than that of 128×128, mainly due to the finer distribution and stronger redundancy of watermarks brought by higher resolution, thus having higher recovery ability under local damage. In addition, the pre-training weights of some deepfake models (such as Uniface, HifiFace, CSCS and FSRT) only support 256×256 input, and 128×128 images need to be resize to high resolution when used. However, the interpolation process will lead to the loss of image details and edge structure, weaken the attack intensity of forgery, and indirectly improve the recoverability and robustness of the watermark.
IV-F Ablation Study
@@
In order to evaluate the contribution of Joint Sub-bands, Attention Module, Frequency Domain Embedding, and GNN in this method, we conducted ablation experiments on both Tracer and Detector. The experiments were conducted on images with a resolution of 128×128. (a) Adjust the sub-bands usage of Tracer and Detector: In Tracer, watermark embedding was reduced from 1,3,4,6 sub-bands to 1,3, while in Detector, the detection was expanded from 1,3 sub-bands to 1,3,4,6. This allows us to evaluate the influence of joint sub-band information redundancy on watermark recovery and deepfake detection. (b) Remove the attention mechanism in the process of watermark embedding to analyze its influence on visual quality. (c) When the DT-CWT frequency domain embedding is removed, the watermark is directly embedded in the U channel. (d) Removing SC-GNN to evaluate its impact on the visual quality of the watermarked image. TABLE VI and VI present the results of the ablation study.
@@
It can be found that: (a) From Case 1, the joint sub-band significantly affects the robustness and detection ability of the watermark. When Tracer only extracts the watermark in 1 and 3 sub-bands, the BER increases obviously, indicating that reducing sub-bands weakens the stability of the watermark. However, when the Detector is extended to 1, 3, 4 and 6 sub-bands, the BER decreases, and the extra sub-bands enhance the redundancy of the watermark, making its detection more stable under common distortion and deepfake forgery. (b) From Case 2, it can be seen that the PSNR and SSIM of the watermark image decrease after the attention mechanism is removed, which indicates that the attention mechanism has optimized the watermark embedding. Although BER has not changed much, the invisibility and visual quality of the watermark have obviously decreased, which verifies the importance of attention mechanism in optimizing visual quality. (c) From Case 3, it can be seen that the robustness of the watermark is significantly reduced after the frequency domain embedding is removed. Although PSNR and SSIM remain at a high level, BER increases significantly under common distortion and deepfake forgery. In addition,the perfect reconstruction, directional selectivity and approximate translation invariance of DT-CWT make frequency domain embedding more resistant to distortion and forgery, which verifies its key role in improving watermark robustness. (d) From Case 4, it can be seen that after removing SC-GNN, the PSNR and SSIM of the watermarked image both decrease, and the BER also shows a slight increase. This indicates that SC-GNN not only plays an important role in optimizing the visual quality of the watermarked image but also provides auxiliary improvements to robustness. By extracting and modeling the structural features of the image, SC-GNN reduces perceptual artifacts, improves invisibility, and enhances resistance to semantic-level manipulations such as deepfake forgeries, thereby contributing to both visual quality and robustness. Compared with more conventional perceptual (e.g., GAN discriminators), our graph-based method places greater emphasis on modeling structural relationships at the regional level. Similarly, compared with Transformer-based constraints that rely on global self-attention to capture long-range dependencies, our method can more effectively maintain spatial consistency in semantically tampered facial images, thereby resisting geometric distortions and local appearance changes, while ensuring both high visual quality and strong robustness of the watermark.
@@ V CONCLUSION
In this paper, a watermarking framework based on frequency domain embedding and SC-GNN is proposed. The robustness is improved by DT-CWT transform, and the detection ability of deepfake is enhanced by combining sub-band information. In addition, the framework leverages dense connectivity to enhance feature reuse and robustness. SC-GNN further optimizes the visual quality of watermark embedding, reduces perceptual artifacts, and improves the invisibility of watermark by attention mechanism. Experimental results show that this method can stably extract watermark and accurately detect tampering in various distorted and forged scenes. Currently, the embedding is performed in the frequency domain of the chrominance channel, which has not fully considered distortions targeting color information. In the future, we plan to explore color-aware embedding strategies to further enhance robustness under color-related perturbations and real-world forgery scenarios. Future work will further improve its generalization ability and optimize its adaptability to complex distortion and unknown deepfake technology.